1) how do you know how to graph this equation without using a calculator? 2x^2 + 8x^2 = 200
Answer by richwmiller(17219)

Are you sure you copied it correctly?
10x^2=200
x^2=20
x=sqrt(20)
x=2sqrt(5)
maybe it should be
2x^2 + 8x = 200
which is a parabola opening up
2x^2 + 8x - 200=0
2*(x^2+4x-100)=0
1/104 (x+2)^2 = 1
x = 2 (sqrt(26)-1)
x = 2 (sqrt(26)+1)
parabola
focus | (-2, -1663/8)=(-2, -207.875)
vertex | (-2, -208)
semi-axis length | 1/8 = 0.125
focal parameter | 1/4
eccentricity | 1
directrix | y = -1665/8